OpenStudy (anonymous):

Charles Gas Law problem: the temperature inside my refrigerator is about 4 degrees Celsius. If I place a balloon in my fridge that initially has temperature of 22 degrees Celsius and a volume of 0.5 liters, what will be the volume of the balloon when it is fully cooled by my refrigerator? The formula is V1/T1 = V2/T2. I need help on how to solve these kind of problems, pleasseee. :]

OpenStudy (mani_jha):

T1=22 degrees Celsius =273+22 Kelvins =295 Kelvins T2=4 degrees celsius=273+4 Kelvins =277 Kelvins V1=0.5 liters V2=? Now, can you substitute the given values into the formula and find out?

OpenStudy (anonymous):

so, 0.5/295K = V2/277K. I still don't really understand how to do these, because I was absent the days my teacher taught it.

OpenStudy (mani_jha):

In these kind of problems, you just have to identify what the initial volume, initial temperature etc would be. In this problem, the balloon is initially outside the fridge. Which means it initially has a temperature equal to the room temperature(22 in this case) and volume equal to its original volume(0.5 liters). Now, when you put the balloon inside the fridge, you'll get your final conditions. The balloon cools down, and it's temperature will eventually become equal to the fridge's temperature(4), and also you'll notice that the balloon contracts(That is Charles' law-Volume decreases when temperature decreases). Now, ok? You just have to think like this.

OpenStudy (anonymous):

Thank you so much :D

OpenStudy (mani_jha):

Glad to help you :)